OpenStudy (anonymous):

f(x) = 3(cos(x))^2 - 6sin(x) 0 ≤ x ≤ 2 Find the intervals on which f is concave down. (Enter the intervals that contain smaller numbers first.) (hint: There are 3 intervals)

myininaya (myininaya):

what did you get for f''

OpenStudy (anonymous):

i got a rock

myininaya (myininaya):

\[f'(x)=3*2*cosx*(-sinx)-6cosx=-6cosxsinx-6cosx=-6cosx(sinx+1)\] \[f''(x)=-6*[-sinx*(sinx+1)+cosx(cosx)]=-6[\cos^2x-\sin^2x-sinx]\] \[f''(x)=-6[1-\sin^2x-\sin^2x-sinx]=-6[-2\sin^2x-sinx+1]\] set f''=0 so we have \[-6[-2\sin^2x-sinx+1]=0\] \[2\sin^2x+sinx-1=0\] \[2\sin^2x+2sinx-sinx-1=0\] \[2sinx(sinx+1)-1(sinx+1)=0\] \[(sinx+1)(2sinx-1)=0\]

myininaya (myininaya):

my f' got cut off

OpenStudy (anonymous):

plus i got 2111 medals, but did anyone congratulate me? \[\color{blue}{\text{noooooo!}}\]

myininaya (myininaya):

congrats satellite

OpenStudy (anonymous):

yeah my \[\color{red}{f'}\] got shot off in the war

myininaya (myininaya):

lol

OpenStudy (anonymous):

too little too late. plus i am smokin'! smoking sensei that is . smoking sensei, feeling irie...

OpenStudy (anonymous):

so what are the intervals?

OpenStudy (anonymous):

are you guys there?

myininaya (myininaya):

you don't see any of the work i did above?

myininaya (myininaya):

sinx+1=0 2sinx-1=0 sinx=-1 sinx=1/2 x=3pi/2 x=pi/6, 5pi/6 |------|--------|-----------|----------| 0 pi/6 5pi/6 3pi/2 2pi plug in a number between each of these numbers to see if is concave up/down use f'' f'' tells you if the function is concave up or concave down

myininaya (myininaya):

plug in a number btw 0 and pi/6 into f''

myininaya (myininaya):

then plug in a number btw pi/6 and 5pi/6 into f'' plug in a number btw 5pi/6 and 3pi/2 into f'' plug in a number btw 3pi/2 and 2pi into f''

myininaya (myininaya):

i do want to help you k? i just don't wanna be rushed

myininaya (myininaya):

f''(.5)=- so concave down on (0,pi/6) f''(1)=+ so concave up on (pi/6,5pi/6) f''(3)=- so concave down on (5pi/6, 3pi/2) f''(5)=- so concave down on (3pi/2, 2pi)

myininaya (myininaya):

do you got it?

OpenStudy (anonymous):

thanks...i ran out of time already!

myininaya (myininaya):

well that sucks are you doing an internet course?